### v1

@@ 5.2 Partition Rounding
Remark 1.
As discussed in Section 2, Miller et al. (2005) have proposed the rounding idea to address risk-averse agents. Our results in this section formally validate that this idea can be extended beyond risk-averse agents to any agents with increasing utility functions. Furthermore, we compare various rounding reductions based on sensitivity, suggest partition rounding reduction for its high sensitivity, and warn that questions cannot be repeatedly used.
5.2.1 Sensitivity of the Partition-Rounded Mechanisms
@@
The CA mechanism requires two questions to compute an individual score, meaning that $K=n/2$. By 4.3 and Lemma 5.3, when the signal space is binary, scoring Alice using $S^{CA}=\sum_{i\in[K]}S^{CA}_{i}$ without rounding is SD-truthful. However, when $|\Sigma|>2$, we have to use the partition rounding reduction of CA to obtain SD-truthful where $S_{\inf}=-1$ and $S_{\sup}=1$. On the bonus question, the probability of receiving a score of 1 is given by $\sum_{\sigma,\sigma^{\prime}}\left(e\cdot J_{\sigma,\sigma^{\prime}}+(1-e)\cdot M_{\sigma,\sigma^{\prime}}\right)T_{\Delta}(\sigma,\sigma^{\prime})$. On the penalty question, the probability of penalizing a score of 1 is given by $\sum_{\sigma,\sigma^{\prime}}\left(e\cdot M_{\sigma,\sigma^{\prime}}+(1-e)\cdot M_{\sigma,\sigma^{\prime}}\right)T_{\Delta}(\sigma,\sigma^{\prime})$. Combining these,
@@
where $(x)^{+}=x$ if $x>0$ and 0 otherwise. Putting these parameters into Eq. 2 returns the sensitivity of the partition rounding reduction for CA. However, note that by 4.3, we do not need to round the individual CA scores when $|\Sigma|=2$. This motivates a variant of an SD-truthful implementation of CA that may have a slightly larger sensitivity in the binary-signal setting. We will introduce and test this implementation in Section 7.
We summarize the limitations of the above three SD-truthful mechanisms. These limitations motivate the search for novel mechanisms, which we explore in the next section.
  • OA does not require the partition rounding reduction to achieve SD-truthfulness and thus has a high sensitivity. However, OA is SD-truthful only when signals are self-dominating.
@@
  • CA uses two questions to compute an individual score, so $K=n/2$. Furthermore, only bonus questions contribute to sensitivity, while penalty questions are used to enforce truthfulness. Therefore, even without rounding, e.g. when $|\Sigma|=2$, only $1/4$ of the questions count for the sensitivity. Moreover, the rounding process divides each score by a factor of 2, which further harms the sensitivity of the mechanism.
6 Enforced Agreement Mechanism
Our discussion highlights the need for more sensitive SD-truthful mechanisms. Specifically, high sensitivity favors a larger number of partitions $K$ and a smaller score range $S_{\sup}-S_{\inf}$. In this section, we ask: can we design mechanisms that use just one question per individual score (so $K=n$) and avoid the need for rounding altogether?
We propose a novel enforced agreement (EA) mechanism which meets the above criteria. We first present EA in the binary-signal setting and then establish a negative result explaining why this approach fails to ensure SD-truthfulness when signals are non-binary. Nevertheless, we show that EA remains to be truthful in non-binary settings when the signal structure is known.
6.1 Binary-Signal Space
The main idea of EA is to control the marginal distribution of Alice’s reports. In the binary setting, the mechanism will commit to a pre-determined empirical distribution $\Phi=(n_{0},n_{1})$ such that $n_{0},n_{1}\in\{0,\ldots,n\}$ and $n_{0}+n_{1}=n$. The mechanism expects Alice to report signal $i$ on exactly $n_{i}$ out of $n$ questions. If Alice reports a different empirical distribution $\Phi_{X_{a}}=(m_{0},m_{1})$ and suppose W.L.O.G. that $m_{0}>n_{0}$, the mechanism will randomly select $m_{0}-n_{0}$ questions on which Alice reports 0 and flip her answers to 1. After enforcing Alice’s empirical marginal distribution to be $\Phi$, the mechanism then scores Alice according to the output agreement mechanism (OA).
Compared with OA, the enforcement process guarantees that Alice will not benefit from over-reporting the majority signal. However, because of the enforcement process, the binary agreement score on each question is no longer i.i.d., meaning that Lemma 5.3 cannot be applied to guarantee SD-truthfulness. Our main result here shows that the EA mechanism is SD-truthful when the information structure satisfies self-prediction (Definition 5).
Theorem 6.1.
We defer the detailed proof to the appendix and present the main ideas below. Note that the final score of the enforced agreement mechanism is the average of $n$ Bernoulli variables each with a potentially different success probability. In the binary-signal setting, there are four possible success probabilities in total — $p_{ij}=\Pr(X_{b}=j\mid X_{a}=i)$ for $i,j\in\{0,1\}$ — representing the probability of Bob observing (and reporting) $j$ on a question conditioned on Alice observing $i$. A Bernoulli variable with success probability $p_{ij}$ corresponds to a question on which Alice observes $i$ but is flipped to $j$ either by the mechanism or by Alice herself. Suppose W.L.O.G. that $m_{0}>n_{0}$, meaning that the mechanism will randomly flip $m_{0}-n_{0}$ reports of Alice from 00 to $1$ under truth-telling. In this case, the final score of Alice is the average of three binomials:
$n\cdot S^{EA}(\tau)\sim{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathrm{Bin}(n_{0},p_{00})}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Bin}(m_{1},p_{11})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathrm{Bin}(m_{0}-n_{0},p_{01})}.$ (3)
$n\cdot S^{EA}(\theta){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sim\mathrm{Bin}(n_{0}-k,p_{00})}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathrm{Bin}(m_{1}-k,p_{11})}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathrm{Bin}(m_{0}-n_{0}+k,p_{01})}+{\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\mathrm{Bin}(k,p_{10})},$ (4)
6.2 Arbitrary Signal Space
Unfortunately, the above idea does not generalize beyond the binary-signal setting. For a general setting with $|\Sigma|=c$, the enforced agreement mechanism can be characterized by an enforced marginal distribution $\Phi=(n_{0},\ldots,n_{c-1})$ and an enforcement rule that ensures an agent’s reports to follow $\Phi$.

### v2

@@ 5.2 Partition Rounding
In Appendix B, we present the sensitivities of the partition-rounded versions of the three mechanisms discussed. We summarize the limitations of these mechanisms below which motivates the design of a new SD-truthful mechanism introduced in the next section.
  • OA is already SD-truthful without rounding and thus has a high sensitivity. However, OA is SD-truthful only for self-dominating signals.
  • PTS requires rounding to ensure SD-truthfulness. Therefore, when the prior of signals is biased, $S^{PTS}_{\sup}$ becomes large, which in turn leads to a low sensitivity.
  • CA uses two questions to compute an individual score, so $K=n/2$. Furthermore, only bonus questions contribute to sensitivity, while penalty questions are used to enforce truthfulness. This dilutes the sensitivity by half again. Therefore, only $1/4$ of the questions count for the sensitivity of the partition-rounded CA. Hence, even without rounding, the sensitivity of the partition-rounded CA is only half of that of the original implementation.
6 The Enforced Agreement Mechanism
This section aims to present a novel mechanism with high sensitivity, which does not require the above rounding reduction. To motivate our idea, we begin by summarizing the key lessons learned from the previous discussions.
First, Example 4.1 shows a particular type of strategy that prevents previous mechanisms from being SD-truthful: reducing/changing the variance of scores, e.g. by always reporting the same signal. This inspires our enforced agreement (EA) mechanism, where the idea is to control a randomness level within agents’ reports so that strategic manipulations cannot arbitrarily alter the variance of the scores.
Second, our analysis of sensitivity in Section 5.2 suggests that a high-sensitivity mechanism requires a larger number of partitions $K$ and a smaller score range $S_{\sup}-S_{\inf}$. This insight guides us to use just one question to compute an individual score (so $K=n$) and avoid the need for rounding. We show that EA meets these criteria.
We first prove the SD-truthfulness of EA in the binary-signal setting, and then establish a negative result explaining why it fails to ensure SD-truthfulness when signals are non-binary. Nonetheless, we show that EA remains truthful in non-binary settings when the signal structure is known or can be accurately learned.
6.1 EA in the Binary-signal Setting
We introduce EA in the binary-signal setting and defer the discussion for general settings to the appendix. The idea of EA is to control the marginal distribution of Alice’s reports. In the binary setting, the mechanism commits to a target empirical distribution $\Phi=(n_{0},n_{1})$ where $n_{0}+n_{1}=n$. The mechanism expects Alice to report signal $i$ on exactly $n_{i}$ out of $n$ questions. If Alice’s actual distribution $\Phi_{X_{a}}=(m_{0},m_{1})$ and suppose W.L.O.G. that $m_{0}>n_{0}$, the mechanism randomly selects $m_{0}-n_{0}$ of her 0-reports and flips them to 1. After enforcing $\Phi$, Alice is scored using the output agreement mechanism (OA).
Compared with OA, the enforcement process in EA prevents Alice from benefiting by over-reporting the majority signal. However, this enforcement introduces dependencies across questions, so the binary agreement scores are no longer i.i.d., and Lemma 5.3 does not apply. Our main result shows that EA is SD-truthful when signals are self-prediction (Definition 4.3).
Theorem 6.1.
[…]
$n\cdot S^{EA}(\tau)\sim{\mathrm{Bin}(n_{0},p_{00})}+{\mathrm{Bin}(m_{1},p_{11})}+{\mathrm{Bin}(m_{0}-n_{0},p_{01})}.$ (2)
$n\cdot S^{EA}(\theta){\sim\mathrm{Bin}(n_{0}-k,p_{00})}+{\mathrm{Bin}(m_{1}-k,p_{11})}+{\mathrm{Bin}(m_{0}-n_{0}+k,p_{01})}+{\mathrm{Bin}(k,p_{10})},$
Additional results are provided in Appendix C. First, we show that EA is not generally SD-truthful in the non-binary setting. Strategic signal permutations can produce score distributions that are not stochastically dominated by truth-telling. However, EA remains truthful when the information structure is known or well estimated. Moreover, we derive a closed-form expression for the sensitivity-maximizing $\Phi$, enabling computation of the optimal enforcement from the information structure.
7 Empirical Evaluations of SD-Truthful Mechanisms
In this section, we empirically compare the sensitivity of the proposed mechanisms across various settings. The results support our theoretical findings: EA consistently exhibits the highest sensitivity among SD-truthful mechanisms in the binary-signal setting (except when the signal prior is nearly uniform). To further demonstrate the practical advantages of EA in effort elicitation, we evaluate the budget efficiency of each mechanism — defined as the minimum payment required to induce a target effort level. Again, EA proves its superiority, which has the best performance against any other SD-truthful and non-SD-truthful mechanisms. This supports our sensitivity analysis, demonstrating that a mechanism with high sensitivity can indeed lead to improved practical performance.
7.1 Datasets and Experiment Setup
We use two real-world datasets to estimate the information structure between two agents. The first dataset contains binary labels classifying whether a compound is appropriate or inappropriate to be synthesized (Baba et al., 2018). The second dataset collects the annotations of the sentiment of $300$ tweets, where the size of the signal space is 4 (Venanzi et al., 2015). We denote these two information structures as $J_{1}$ and $J_{2}$ respectively, and defer the details to Section E.1.
In the main body of the paper, we focus on SD-truthful mechanisms, which include:
  • OA (Section 4.1), which is SD-truthful only when signals are self-dominating.
  • PTS-partition-round—the partition rounding reduction of PTS (Section 4.2), where we re-weight every individual score using signal prior and take the average.
  • CA-partition-round—the partition rounding reduction of CA (Section 4.3), where we create $K=n/2$ partitions of questions and score agents the average of the individual CA score for each partition after rounding.
  • MA-partition-round—the partition rounding reduction of the matching agreement mechanism (Appendix D).
  • EA-prior—EA with enforcement $\Phi=n\cdot\Pr(X_{a})$.
  • EA-uniform—EA with $\Phi_{i}=n\cdot 1/|\Sigma|,~\forall i\in\Sigma$.
  • EA-optimal—EA with the sensitivity-maximizing $\Phi$ computed according to Section C.3.
For information structures with binary-signal settings, sensitivities can be computed analytically as shown in Appendix B. For general cases, we estimate sensitivity using a Monte Carlo approach. In particular, we simulate reports to each of the $n$ questions with both agents exerting full effort $e=1$ and run $\mathcal{M}$ to compute the scores. Repeating this process for T=20, [^0] T=20,\!000 times yields $T$ i.i.d. samples of $S^{\mathcal{M}}(e)$. We then compute the score of Alice when she deviates to a lower effort level at $e-\Delta e=0.8$, and obtain $T$ samples of $S^{\mathcal{M}}(e-\Delta e)$. The sensitivity of at effort $e$ is then estimated as $\frac{\mathbb{E}[S^{\mathcal{M}}(e)]-\mathbb{E}[S^{\mathcal{M}}(e-\Delta e)]}{\Delta e}\cdot\frac{1}{\operatorname{std}(S^{\mathcal{M}}(e))}$.
7.2 The Sensitivity of SD-Truthful Mechanisms
We now compare the sensitivity of the discussed SD-truthful mechanisms. In Figure 1(a), we show the sensitivity when the prior of the binary signal is varied, i.e. $\Pr[X_{a}=0]\in[0.1,0.9]$, and the number of questions $n=100$.
Figure 1(b) and 1(c) present sensitivity as the number of questions $n\in\{10,20,\ldots,100\}$ increases, under information structures $J_{1}$ and $J_{2}$, respectively. In $J_{2}$, where $|\Sigma|>2$, EA is no longer SD-truthful; however, we still include its sensitivity for comparison.
(a) Varying prior $\Pr[X_{a}=0]$
(b) Varying #tasks under $J_{1}$
(c) Varying #tasks under $J_{2}$
Figure 1: A Comparison of the Sensitivity of Different Mechanisms.
We make the following observations. First, although the partition reductions of CA and MA are SD-truthful for general settings, their sensitivities are dominated by other SD-truthful mechanisms. This is due to their requirements of the penalty questions, which do not contribute to the sensitivity, and the fact that they require two questions to compute an individual score.
Second, PTS performs well when the signal prior is nearly uniform, i.e. $\Pr(X_{a}=0)\approx 0.5$. Intuitively, this is because the rounding factor that enlarges the gap $S^{PTS}_{\sup}-S^{PTS}_{\inf}$ shrinks when the prior is close to uniform. However, even a slight bias, i.e. $|\Pr(X_{a}=0)-0.5|>0.03$ causes PTS-partition-round to be outperformed by EA.
Third, we observe that enforcing the signal prior in EA closely approximates the optimal enforcement. Remarkably, even EA-uniform, a detail-free version that doesn’t require knowledge of the prior, consistently outperforms PTS-partition-round, which does rely on prior knowledge for SD-truthfulness.
Additionally, we note that the partition-rounding reduction of MA coincides with that of CA when signals are binary (see Appendix D). Consequently, the blue and red overlap in Figure 1(a) and 1(b).
7.3 Budgetary Efficiency
We further evaluate the performance of the proposed mechanisms in an effort-elicitation setting, aiming to demonstrate that SD-truthful mechanisms with higher sensitivity lead to more efficient effort elicitation. This setting differs from the sensitivity estimation setup in several key ways. First, it involves more than two agents, with each agent potentially answering different subsets of questions. Second, we consider a symmetric effort equilibrium in which all agents exert the same effort and no agent has an incentive to deviate. In contrast, in the previous setting, Bob was assumed to always exert full effort. Finally, we incorporate effort costs and explicit payment schemes to better reflect realistic crowdsourcing conditions. These enhancements are intended to strengthen the robustness of our findings.
7.3.1 Model
Following a prior work (Zhang and Schoenebeck, 2023a), we consider an effort-elicitation model where each agent has an effort level $e_{i}$ with a convex cost function $c(e_{i})=e_{i}^{2}$ reflecting the increasing marginal cost of effort. The principal pays each agent a function of her score computed by a peer prediction mechanism, and agents best respond by choosing an effort that maximizes their expected utility — the difference between the expected payment and the cost of effort. Note that because the peer prediction score of an agent depends on all agents’ effort levels, we consider a symmetric equilibrium as our solution concept, where every agent exerts the same effort level, and no one wants to unilaterally deviate. This motivates the use of budgetary efficiency as a performance metric for peer prediction mechanisms, defined as the minimum budget required to incentivize a target effort level at the symmetric equilibrium. Intuitively, a mechanism that is more sensitive to changes in information can induce higher effort under the same budget and therefore achieve greater budgetary efficiency.
  • The linear payment scheme compensates agent $i$ according to
    $\pi_{i}=a\cdot S^{\mathcal{M}}_{i}+b.$
  • The nonlinear threshold payment scheme awards agent $i$ a fixed bonus if her score exceeds a threshold $t$:
    $\pi_{i}=C_{\operatorname{thres}}\cdot\mathbbm{1}[S^{\mathcal{M}}_{i}>t].$
    In our experiments, we first estimate the score distribution by simulating agents’ reports and applying the peer prediction mechanism. We then set the threshold $t$ as a chosen percentile of the resulting empirical score distribution, and keep it fixed when computing the final payments.
In the above payment schemes, $a,b,C_{\operatorname{thres}},t$ are constant parameters that can be optimized for budgetary efficiency in practice, given the score distributions. Additionally, we require payments to satisfy limited liability (i.e. the ex-post payment must be non-negative) and individual rationality (IR, i.e. the expected utility must be non-negative).
Note that both the peer prediction mechanism and the associated payment scheme influence budgetary efficiency. To preserve truthfulness in payments, a truthful mechanism must be paired with a linear payment scheme. In contrast, SD-truthful mechanisms can be combined with non-linear payment schemes, such as threshold payments. Our goal is to investigate whether the combination of SD-truthful mechanisms and non-linear payments can achieve greater efficiency than truthful mechanisms with linear payments.
7.3.2 Experiment setup
We consider the information structure $J_{1}$ with 50 agents and 500 tasks, where each agent works on 50 tasks and each task is randomly assigned to 5 agents. To incentivize a desired effort level $e$ as a local symmetric equilibrium, we learn the parameters in the payment schemes by solving $\nabla\pi_{i}(e_{i})=\nabla c(e_{i})$ for any $i$. Note that $\nabla c(e_{i})$ can be directly computed given a desired effort $e_{i}=e$. To compute $\nabla\pi_{i}$, we repeat the same procedure as discussed in Section 7.1. In particular, we simulate samples of signals and apply various mechanisms to compute the peer prediction score $S_{i}^{\mathcal{M}}$ for each agent. Then, varying the effort from $e$ to $e-\Delta e$, we can estimate $\nabla S^{M}_{i}$ using a sample size T=10, [^0] T=10,\!000. Putting this estimation into the formula of various payment schemes thus gives us an estimation of $\nabla\pi_{i}$. After learning the parameters of the payment schemes, we compute the budgetary efficiency as the total payment to incentivize such an effort level. For the threshold payment, we split the T=10, [^0] T=10,\!000 samples into two halves: the first half is used to compute the threshold $t$ as the 75th percentile of the score distribution across all agents, and the second half is used to estimate the resulting payment.
(a) $e=0.6$
(b) $e=0.7$
(c) $e=0.8$
(d) $e=0.9$
Figure 2: Estimated Sensitivity v.s. Total Payment under Threshold Payment.
7.3.3 Results
We compute the optimal linear payments — subject to limited liability and individual rationality — for all truthful mechanisms introduced in Section 7.1, and compare them against the threshold payments used with SD-truthful mechanisms. A lower total payment indicates greater budgetary efficiency. Figure 2 presents the results for SD-truthful mechanisms combined with threshold payments across four target effort levels. The red dashed line marks the lowest total payment required by any truthful mechanism under a linear payment scheme, while the blue dashed line indicates the minimum payment required to satisfy individual rationality. Each dot represents the total payment to elicit the desired effort for a given mechanism, ignoring the IR constraint.
We make several key observations. First, sensitivity correlates strongly with budgetary efficiency, consistent with the theoretical findings of prior work (Xu et al., 2024; Zhang and Schoenebeck, 2023a). Second, as predicted by our theoretical insights, the EA mechanism — particularly when enforcing the signal prior — consistently achieves the best performance. Third, the advantage of pairing an SD-truthful mechanism with a threshold payment scheme over a truthful mechanism with a linear payment scheme becomes more pronounced at higher target effort levels. This is because greater effort leads to more accurate peer prediction scores, thereby increasing sensitivity, which non-linear payment can leverage more effectively. These results highlight the practical importance of designing highly sensitive SD-truthful mechanisms, as they offer substantial improvements in the performance of incentive systems across real-world settings.
8 Conclusion and Future Work
This paper initiates the discussions of stochastically dominant-truthful (SD-truthful) peer prediction mechanisms, which ensure that in equilibrium, the score distribution under truth-telling first-order stochastically dominates any other strategy. Unlike traditional peer prediction mechanisms, which rely on linear utility assumptions, SD-truthful mechanisms can preserve truthfulness for any increasing utility function. We show that existing mechanisms fail to naturally achieve SD-truthfulness and propose a rounding method to enforce it. However, this process often reduces the sensitivity of the original mechanism, making it less efficient in practice. Additionally, we introduce the Enforcement Agreement (EA) mechanism, which is empirically shown to have the highest sensitivity among all SD-truthful mechanisms in the binary-signal setting—except when the signal prior is nearly uniform.
Future work is needed to investigate more sensitive SD-truthful mechanisms beyond the binary-signal setting. A promising next step is to explore approximate SD-truthfulness, which allows truth-telling to lose a small utility for some utility functions. This relaxation could lead to mechanisms with better sensitivity. Furthermore, SD-truthfulness alone does not ensure that truth-telling maximizes expected rewards in a tournament (rank-order rewards). Because agents’ scores are correlated, a manipulation strategy with a dominated score distribution may still raise an agent’s chance of ranking above others by altering these correlations. An important direction for future work is therefore the design of peer prediction mechanisms that provide incentive guarantees under tournaments.
Broader Impact
On the positive side, our method enhances the reliability and efficiency of crowdsourced data collection, which supports more trustworthy machine learning systems and reduces reliance on expert annotation. However, mechanisms that incentivize effort may unintentionally disadvantage contributors with limited time, resources, or familiarity with the task, raising fairness concerns. We aim to mitigate these risks by designing robust, broadly applicable mechanisms and promoting transparency in their deployment.
Acknowledgment
This work was partly carried out at the DIMACS Center at Rutgers University and supported by NSF #2313137.
Appendix A Continued Related Work: Mutual Information Mechanisms
We review several classic peer prediction mechanisms that evaluate agents by estimating mutual information, and illustrate why these mechanisms fail to achieve SD-truthfulness. First, proposed by Kong and Schoenebeck [2019], we can estimate the empirical point-wise joint distribution between $\hat{X}_{a,\cdot}$ and $\hat{X}_{b,\cdot}$, using pairs of reports from two agents, $(\hat{X}_{a,j},\hat{X}_{b,j})_{j\in[n]}$. Note that because questions are i.i.d., we can use reports from all questions to estimate the same distribution. Then, plugging the learned empirical joint distribution into an $f$-divergence operator returns a (biased) estimate of the $f$ point-wise mutual information. The following is the score for the $f$-mutual information mechanism [Kong and Schoenebeck, 2019]:
$S^{fMI}=D_{f}\left(\Pr\left(\hat{X}_{a,\cdot},\hat{X}_{b,\cdot}\right)\|\Pr\left(\hat{X}_{a,\cdot}\right)\Pr\left(\hat{X}_{b,\cdot}\right)\right).$
However, because the estimation is biased, the $f$-mutual information mechanism is only asymptotically truthful, meaning that it requires an infinite number of i.i.d. questions to satisfy Definition 3.1. When the number of tasks is small, there is a non-zero probability that truth-telling is not the best response. For example, if Alice observes “yes” on all her questions, which happens with a positive probability, then she knows that truth-telling will always result in a score of 0 regardless of Bob’s responses. This is because the estimated point-wise joint distribution is the same as the product of marginal distributions. However, by misreporting “no” on some questions, she has a non-zero probability of receiving a positive score. In comparison, such deviations are discouraged under the enforced agreement (EA) mechanism because the mechanism will randomly flip the responses on behalf of the agent (see Section 6).
Several follow-up works, such as the Determinant Mutual Information (DMI) mechanism by Kong [2020] and the pairing mechanism by Schoenebeck and Yu [2020], propose different ideas to estimate the mutual information. However, they are still not SD-truthful. Consider the strategy from Example 4.1, where Alice always reports “yes.” Under both DMI and the pairing mechanism, this strategy always yields a score of zero. In contrast, similar to the CA mechanism, there is a positive probability that truth-telling may result in a negative score. Thus, this constant-reporting strategy is not dominated by truth-telling under these mutual information-based mechanisms whose score can take negative values.
Appendix B Sensitivity of Prior Mechanisms Under Partition Rounding
@@
The CA mechanism requires two questions to compute an individual score, meaning that $K=n/2$. By F.1 and Lemma 5.3, when the signal space is binary, scoring Alice using $S^{CA}=\sum_{i\in[K]}S^{CA}_{i}$ without rounding is SD-truthful. However, when $|\Sigma|>2$, we have to use the partition rounding reduction of CA to obtain SD-truthful where $S_{\inf}=-1$ and $S_{\sup}=1$. On the bonus question, the probability of receiving a score of 1 is given by $\sum_{\sigma,\sigma^{\prime}}\left(e\cdot J_{\sigma,\sigma^{\prime}}+(1-e)\cdot M_{\sigma,\sigma^{\prime}}\right)T_{\Delta}(\sigma,\sigma^{\prime})$. On the penalty question, the probability of penalizing a score of 1 is given by $\sum_{\sigma,\sigma^{\prime}}\left(e\cdot M_{\sigma,\sigma^{\prime}}+(1-e)\cdot M_{\sigma,\sigma^{\prime}}\right)T_{\Delta}(\sigma,\sigma^{\prime})$. Combining these,
@@
where $(x)^{+}=x$ if $x>0$ and 0 otherwise. Putting these parameters into Equation 1 returns the sensitivity of the partition rounding reduction for CA. However, note that by F.1, we do not need to round the individual CA scores when $|\Sigma|=2$. This motivates a variant of an SD-truthful implementation of CA that may have a slightly larger sensitivity in the binary-signal setting. We will introduce and test this implementation in Section 7.
We summarize the limitations of the above three SD-truthful mechanisms.
  • OA does not require the partition rounding reduction to achieve SD-truthfulness and thus is likely to have a high sensitivity. However, OA is SD-truthful only when signals are self-dominating.
@@
  • For the CA mechanism, only bonus questions contribute to the sensitivity of the mechanism, while penalty questions are used to enforce truthfulness. However, CA requires two questions to compute a single individual score, which means K=n/2K=n/2.333We emphasize that increasing $K$ by rerunning partitions is not SD-truthful (Example 4.1). Therefore, even without rounding, e.g. when $|\Sigma|=2$, the number of effective questions is only one half of the available questions. Moreover, the rounding process divides each score by a factor of 2, which further harms the sensitivity of the mechanism.
Appendix C The Enforced Agreement Mechanism (Continued)
C.1 Arbitrary Signal Space
Unfortunately, we show that the above idea does not generalize beyond the binary-signal setting. For a general setting with $|\Sigma|=c$, the enforced agreement mechanism can be characterized by an enforced marginal distribution $\Phi=(n_{0},\ldots,n_{c-1})$ and an enforcement rule that ensures an agent’s reports to follow $\Phi$.